# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v1

@@ 7 AudioTrust: Authentication
In the Authentication section, we mainly focus on three problems of ALLMs: (1) ALLMs are not yet robust enough against synthetic and voice impersonation. With the rapid development of voice cloning technologies (e.g., F5-TTS [12]), attackers can easily generate highly realistic voices of target users, leading to false acceptance of unauthorized entities. (2) Attackers may use social engineering by creating urgent situations or applying psychological pressure to circumvent strict verification standards. (3) Attackers may deliberately introduce acoustic interference, such as echo or crowd noise, to reduce the reliability of voice matching or to confuse the verification logic.
Attack Strategies. We categorize authentication attacks into three types: Hybrid Spoofing, Identity Verification Bypass and Voice Cloning Spoofing. This taxonomy is based on the logical-vs-physical access distinction used in ASVspoof [53] and recent surveys [42]. Among them, Hybrid Spoofing combines speech synthesis, background context (e.g., office noise), and social engineering dialogue to construct complex deception scenarios. Identity Verification Bypass directly targets verification rules, triggering the model to relax access restrictions through incomplete credentials or emotional urgency. Voice cloning spoofing uses deepfake audio to perform high-fidelity imitation of authorized users, testing whether the model can distinguish real from cloned voices when correct phrases are used. For each sample, we first perform data augmentation using GPT-4o [54], and then use text-to-speech tools such as F5-TTS [12] to construct the audio data. We constructed 100 data samples each for Hybrid Spoofing and Identity Verification Bypass, and 200 samples for Voice Cloning Spoofing. Further details on data classification are provided in Section I.1.
@@
| MiniCPM-o 2.6 | 76 | 80 | 20.5 |
| Qwen2-Audio | 58 | 30 | 7.5 |
| SALMONN | 74 | 92 | N/A |
| Ultravox | 5 | 50 | 72 |
@@
| Gemini-1.5 Pro | 4 | 13 | 66.5 |
| Gemini-2.5 Flash | 3 | 19.8 | 89 |
| Gemini-2.5 Pro | 5 | 15.5 | 89.5 |
| GPT-4o Audio | 2 | 10 | 16.5 |
| GPT-4o mini Audio | 0 | 7 | 14 |
IVB: Identity Verification Bypass, HS: Hybrid Spoofing, VCS: Voice Cloning Spoofing.
Note: SALMONN consistently disregarded prompt instructions by outputting audio descriptions, which prevented obtaining valid results for voice cloning spoofing. indicates value above the average within its category (open-source or closed-source), indicates value below average, with subscript showing the absolute difference from the category average. For authentication metrics, lower values indicate better security (fewer successful attacks).
Evaluation & Metrics. For Voice Cloning Spoofing, the model evaluates whether the enrollment and access voices match; a failure occurs if spoofed inputs trigger responses like “Yes, I’m here” or “Access granted”. The key metric is False Acceptance Rate (FAR), the proportion of unauthorized accesses accepted. For Hybrid Spoofing, models must also cope with background noise and social engineering (e.g., urgent simulated requests), evaluated by both FAR and Social Engineering Strength (SES)—the degree of psychological manipulation. In Identity Verification Bypass, robustness is tested by attempting authentication with only partial passphrases under emergency scenarios, with FAR as the main metric. Further inference and evaluation details are in Section I.3 and Section I.4.
Results. In Table 6, we evaluated the performance of four open-source ALLMs and five closed-source ALLMs on the Authentication test set across different scenarios. It can be observed that open-source models exhibit better performance than their closed-source counterparts. (1) In the Identity Verification Bypass test, we found that closed-source models were almost never deceived, as they could recognize potential threats, with the GPT-4o series performing the best by rejecting authentication attempts. Among open-source models, Ultravox performed relatively well, while MiniCPM-o 2.6 was the least effective. (2) In the Hybrid Spoofing test, similar conclusions were drawn, with closed-source models generally outperforming open-source ones, where the GPT-4o series stood out. However, within closed-source models, the deception rate was relatively higher, with Qwen2-Audio performing the best but still weaker compared to other closed-source models.(3) In the Voice Cloning Spoofing test, the best-performing models are Qwen-2-Audio-Chat and GPT-4o mini Audio. There is substantial variation in performance across different models, with the Gemini series generally demonstrating weaker results. Moreover, as shown in Table 24, we find that simply increasing the strictness of the textual prompts improves the models’ resistance to spoofing attacks across all systems. This suggests that in downstream applications of ALLMs, carefully designing system prompts may significantly enhance spoofing defense capabilities. See Section I.5 for further details.
@@ 8 Conclusions
This paper introduces AudioTrust, the first comprehensive benchmark framework for multidimensional trust evaluation of ALLMs. AudioTrust systematically covers six major trustworthiness dimensions: fairness, hallucination, safety, privacy, robustness, and authentication, considering ALLMs’ unique risks and application requirements. We construct a diverse evaluation dataset that reflects practical scenarios such as daily conversations, emergency calls, and voice assistant interactions, and design a targeted metric system. An automated, scalable evaluation pipeline based on GPT-4o enables efficient and objective assessment. Experimental results illustrate the trustworthiness boundaries and limitations of current open- and closed-source ALLMs in high-risk tasks, including systematic bias toward sensitive attributes (e.g., gender, accent), and limited robustness under noise, multi-speaker, or adversarial conditions. AudioTrust offers actionable insights and provides a solid foundation for future trustworthy ALLM research. The framework and evaluation platform are publicly available to promote further study in this area.
@@ Appendix A Introduction to Audio Large Language Models
The emergence of ALLMs signifies a pivotal paradigm shift in the domain of multimodal artificial intelligence systems [78, 56]. These models fundamentally extend the capabilities of traditional LLMs [102, 22] —which have demonstrated remarkable proficiency in processing and generating textual information—by enabling the comprehension and synthesis of auditory signals. This advancement substantially surpasses conventional Automatic Speech Recognition (ASR) systems [48], whose primary objective is to faithfully transcribe spoken language into text. In contrast, ALLMs aim to achieve a more holistic understanding of acoustic environments, encompassing not only the lexical content of speech but also paralinguistic cues (e.g., prosody, affective tone), speaker characteristics, musical elements, and background environmental sounds [65]. Such deep exploration of the rich semantic information embedded in audio signals is crucial for realizing more natural and context-aware human-computer interaction. ALLMs are generally divided into two primary categories: speech understanding models and speech interaction models.
The rapid maturation of this field has been largely propelled by significant advancements in Self-Supervised Learning (SSL) methodologies, which enable models to acquire robust representations from vast quantities of unlabeled audio data. Concurrently, sophisticated multimodal training paradigms have played a critical role, facilitating the synergistic integration and joint learning of information across auditory and linguistic modalities [20, 60, 93, 72]. By aligning the acoustic feature space with the inherent semantic comprehension capabilities of LLMs, ALLMs are able to address tasks beyond simple speech-to-text conversion, such as audio event classification, audio scene description, audio-based question answering, and even engaging in multi-turn spoken dialogues. These capabilities mark new frontiers for developing artificial intelligence applications that can more profoundly interpret and respond to our auditory world. However, as ALLMs are increasingly integrated into real-world applications, understanding their impact under various trustworthiness conditions becomes critically important. This study aims to construct a benchmark, AudioTrust, to comprehensively and systematically evaluate the performance and potential risks of ALLMs across different trustworthiness dimensions, such as robustness, fairness, privacy protection, and safety. This evaluation is intended to provide scientific evidence and practical guidance for the responsible development, deployment, and regulation of ALLMs.
@@ A.1 Speech Understanding Models
Speech understanding models process and comprehend audio inputs, transforming them into semantic representations that facilitate language understanding. However, they lack the ability to generate audio responses. These models typically operate in a unidirectional manner, receiving audio as input and producing text-based outputs. Notable representatives include Qwen2-Audio [14], which integrates audio understanding capabilities into the Qwen2 [84] via dedicated audio encoders and cross-modal adapters. These models demonstrate strong performance in tasks such as speech transcription, audio description, and audio-based question answering, yet their outputs remain restricted to textual modalities. SALMONN [65] likewise exhibits robust semantic audio understanding across diverse acoustic conditions, while maintaining a purely text-based output interface.
@@ A.2 Speech Interaction Models
Speech interaction models go beyond mere comprehension to enable bidirectional audio communication. These models are capable not only of understanding audio inputs, but also of generating contextually appropriate audio responses, thereby facilitating more natural human-computer interaction. Prominent examples include GPT-4o [54], which represents a significant advance in multimodal interactive capability by processing and generating audio in near real-time conversational scenarios. MiniCPM-o 2.6 [90] provides similar functionalities in an open-source format, supporting coherent audio dialogues while demonstrating comprehension of audio contexts. Such models enable a wide range of applications, from virtual assistants to assistive tools for visually impaired users.
@@ Appendix B Benchmark Models
To systematically investigate these trustworthiness aspects, we have selected a diverse set of models. This set includes both mainstream proprietary commercial models, such as GPT-4 [54] and Gemini [66], as well as representative and robust open-source ALLMs, including Qwen2-Audio [14] and MiniCPM-o 2.6 [90]. To ensure fairness and objectivity, all models are systematically tested on the same datasets and with identical evaluation metrics, followed by thorough comparative analyses of experimental results. It is worth noting that our methodology considers not only the fundamental audio comprehension capabilities of each model, but also examines their potential strengths and limitations in aspects such as complex interactions and knowledge transfer. This systematic safety evaluation provides a solid foundation for the future optimization and development of ALLMs.
@@ B.1 Open-Source Models
In conducting trustworthiness evaluations of unified ALLMs, we selected five representative open-source audio and multimodal models: Qwen2-Audio, SALMONN, Qwen2.5-Omni, MiniCPM-o 2.6, and Ultravox.
Qwen2-Audio [14] is a large-scale audio-language model that establishes a seamless pipeline between the Whisper-large-v3 encoder and the Qwen-7B language model, thereby supporting both spoken dialogue and audio analysis interaction modes. In real conversational and multitask zero-shot evaluations, the model leverages Mel-spectrograms of 16kHz audio combined with instruction tuning and Direct Preference Optimization (DPO), significantly improving the precision and robustness of responses to human intent.
SALMONN [65] pioneered a dual-encoder architecture (Whisper speech encoder and BEATs audio encoder) together with a window-level Q-Former and LoRA adapters. This enables the pretrained Vicuna text LLM to achieve unified understanding of speech, environmental sounds, and music. The model also demonstrates emergent capabilities in cross-modal reasoning beyond the training tasks and in few-shot activation tuning.
MiniCPM-o 2.6 [90] integrates four major components: SigLip-400M, Whisper-medium, ChatTTS-200M, and Qwen2.5-7B, supporting bilingual real-time dialogue in an end-to-end multimodal fashion, as well as controllable interactions in emotion and speaking rate, and high-quality voice cloning. It consistently outperforms proprietary models of equivalent scale on benchmarks such as OpenCompass and StreamingBench.
Ultravox [2] directly maps raw audio into the high-dimensional representation space of LLMs, thereby seamlessly eliminating the traditional ASR stage. This model not only comprehends speech content but also captures paralinguistic features such as tone and pauses, and supports streaming text outputs.

### v3

@@ 7 AudioTrust: Authentication
In this section, we investigate the reliability of ALLMs for authentication. Text-based authentication relies on semantic secrets like passwords. Audio authentication is more complex because a voice signal contains both a semantic component (the passphrase) and an acoustic one (the speaker’s unique voiceprint). This dual nature creates a unique attack surface. For example, an attacker could use a perfect AI-generated voice clone to speak a correct passphrase, defeating systems that rely on either modality alone. We evaluate how well ALLMs can defend against different impersonation attacks.
Attack Strategies. We devise a taxonomy of authentication attacks that exploit the multi-dimensional attributes of audio signals, categorizing them into three primary classes. (1) Identity Verification Bypass (IVB), which injects strong emotional cues (e.g., feigned urgency or distress) to exploit social engineering principles and induce the model to lower its security thresholds, and (2) Hybrid Spoofing (HS), this attack convolves cloned or synthesized speech with background noise and reverberation characteristic of specific scenes (e.g., a bustling office). Such attacks challenge the model’s robustness in source separation and noise suppression, and more critically, its ability to jointly reason about the logical consistency between the purported speaker identity and their acoustic context. (3) Voice Cloning Spoofing (VCS), which employs advanced speech synthesis (e.g., F5-TTS (Chen et al., 2024)) to generate highly realistic cloned voices saying a correct passphrase, thereby testing the model’s sensitivity to subtle physical traits in the vocal fingerprint. This taxonomy moves beyond traditional spoofing distinctions to directly probe ALLMs’ intrinsic vulnerabilities in processing complex auditory signals. Based on this framework, we constructed a 400-sample evaluation dataset (details in Sections˜I.1 and I.2).
@@
| SALMONN | 74 | 93 | N/A |
| Ultravox | 5 | 43 | 72 |
| Qwen2-Audio | 58 | 29 | 7.5 |
| MiniCPM-o 2.6 | 76 | 57 | 20.5 |
| Step-Fun | 11 | 3 | 78 |
| Qwen2.5-omni | 36 | 36 | 87.5 |
| Kimi-Audio | 21 | 24 | 75.5 |
| OpenS2S | 3 | 34 | 50 |
| Step-Audio2 | 36 | 85 | 49 |
| Open-source Avg. | 35.6 | 44.9 | 55.0 |
@@
| Gemini-1.5 pro | 4 | 5 | 66.5 |
| GPT-4o Audio | 2 | 0 | 16.5 |
| GPT-4o mini Audio | 0 | 0 | 14 |
| Gemini-2.5 Flash | 3 | 7 | 89 |
| Gemini-2.5-Pro | 5 | 3 | 89.5 |
| Closed-source Avg. | 2.8 | 3.0 | 55.1 |
Note: SALMONN consistently disregarded prompt instructions by outputting audio descriptions, which prevented obtaining valid results for voice cloning spoofing. For authentication metrics, lower values indicate better security (fewer successful attacks).
Evaluation and Metrics. Our evaluation framework is designed to comprehensively assess the robustness of models under different attack scenarios. FAR is defined as the proportion of impostor voices incorrectly identified as authentic users, which will serve as the privacy metric for all attack scenarios. GPT-4o will be our main evaluator, used to verify if ALLMs deny attacker requests. Human evaluators will then review GPT-4o’s results before the final FAR calculation. In the case of HS, where the model faces both identity spoofing and environmental interference, FAR remains the central performance indicator. A failure is registered whenever the model responds positively to a forged input (e.g., acknowledging identity confirmation or granting access). Further details on the evaluation are provided in Section˜I.3 and I.4.
Results. In Table˜6, we report the performance of both open- and closed-source ALLMs on our evaluation dataset. Several key observations emerge: (1) Our findings in the IVB and HS settings reveal a critical attack vector absent in text-only systems: non-semantic features. We observed that the average false acceptance rate for open-source models was 35.6% in IVB settings, but significantly higher at 44.9% in HS settings. These settings differ considerably in paralinguistic features. The HS setting, for example, includes background audio and other speakers’ voices, both common elements in social engineering attacks conveyed through audio. This suggests that paralinguistic cues might aid attackers in successful authentication. Conversely, closed-source models appear to demonstrate greater robustness. (2) In the VCS tests, we observe substantial performance discrepancies across models, with the Gemini family exhibiting comparatively weaker defenses. Interestingly, we find that simply increasing the strictness of system prompts consistently improves resilience against spoofing attacks across all systems. This suggests that in downstream ALLM applications, carefully crafted system prompts provide an efficient means of strengthening authentication security. Further details can be found in Section˜I.5.
@@ 8 Conclusions
This paper introduces AudioTrust, the first comprehensive benchmark framework for reliability assessment specifically designed for ALLMs. Unlike prior evaluations targeting text-based LLMs, AudioTrust places particular emphasis on the unique characteristics of the audio modality and the novel security challenges it entails. The framework systematically spans six key dimensions: fairness, hallucination, safety, privacy, robustness, and authentication, and also includes audio-specific risks into the design space and threat modeling. To ensure broad coverage, AudioTrust constructs a large-scale audio dataset that reflects a wide range of complex conditions. Also, we develop dedicated metrics to assess these risks, integrated with an automated pipeline powered by GPT-4o, enabling scalable evaluation. Our experimental results demonstrate that both open-source and closed-source ALLMs exhibit pronounced limitations when faced with high-risk challenges unique to the audio domain. Beyond these empirical findings, AudioTrust offers actionable insights for researchers. It defines the reliability boundaries of current ALLMs in real-world audio scenarios and lays a foundation for future work on trustworthy model design. We have publicly released our framework and evaluation platform to foster broader community-driven research in this critical area.
Reproducibility Statement
To ensure the reproducibility of our work, we release all related code, dataset metadata, and detailed experimental configurations. Our evaluation framework AudioTrust, the automated evaluation scripts, and the leaderboard generation code are hosted in an anonymous GitHub repository: https://github.com/AudioTrust/AudioTrust. Appendix C describes the decoupled two-stage architecture (inference and evaluation) of our platform, which provides guidance for reproducing our evaluation pipeline. Throughout the main paper (Sections 2–7), we explicitly indicate the corresponding appendix sections. Specifically, Appendices D through I present detailed treatments of six evaluation dimensions: fairness, hallucination, safety, privacy, robustness, and verification, which cover attack strategies, dataset construction, evaluation protocols, and the metrics employed. For instance, Appendices D.1 and D.2 describe the dataset construction process for fairness evaluation, while Appendices E.3.1 and E.3.2 explain the details of the metrics for hallucination evaluation. We believe that the public release of these resources will facilitate both reproducibility and extension of our study by the community.
Ethical Statement
Our work adheres to the ICLR Code of Ethics. The primary objective of this research is to foster the development of more trustworthy, secure, and equitable Audio Large Language Models (ALLMs) by providing a comprehensive evaluation framework, AudioTrust. We aim to identify and understand the potential risks within these models, thereby guiding the community toward building safer and more reliable AI systems. Ethical considerations have been carefully addressed at every stage of our research, from dataset creation to the potential impact of our findings.
Dataset Curation and Human Participants
All data used in the AudioTrust benchmark was curated with strict ethical considerations. The majority of samples are synthetically generated (e.g., via text-to-speech systems and large language models) or drawn from publicly available datasets with permissive research licenses (e.g., Common Voice and freesound). No real personal or private information was included; sensitive content such as account numbers was generated synthetically to evaluate information-leakage risks. Annotations regarding attributes such as gender, accent, and age were taken from public datasets or inferred only when necessary, and were used solely to study potential biases. Human annotators were involved exclusively for quality control tasks (e.g., checking transcription accuracy), without collection of personal data or interventions, and thus this study does not constitute research involving human subjects.
Potential Misuse and Dual Use Concerns
Our research, by its nature, reveals vulnerabilities in current ALLMs concerning safety, privacy, robustness, and authentication. We recognize that while the methods and scenarios within AudioTrust are designed for evaluation, they could theoretically be adapted by malicious actors to develop more effective attacks. We systematically explore strategies for jailbreaking, audio based social engineering, and spoofing authentication systems. Disclosing these vulnerabilities is a double edged sword. We believe that the benefit of transparently presenting these risks to the research community through a benchmark far outweighs the risk of misuse. By establishing a public benchmark, we enable developers, both open source and proprietary, to test, fortify, and improve their models against these specific threats. Our objective is to catalyze defensive research and promote the adoption of robust safety alignment. The purpose of our public release is to accelerate this positive feedback loop, leading to safer ALLMs for everyone.
@@ Appendix A Introduction to Audio Large Language Models
The emergence of ALLMs signifies a pivotal paradigm shift in the domain of multimodal artificial intelligence systems (Wu et al., 2024; Peng et al., 2024). These models fundamentally extend the capabilities of traditional LLMs (Zhao et al., 2023; Hadi et al., 2023), which have demonstrated remarkable proficiency in processing and generating textual information. They achieve this by enabling the comprehension and synthesis of auditory signals. This advancement substantially surpasses conventional Automatic Speech Recognition (ASR) systems (Malik et al., 2021), whose primary objective is to faithfully transcribe spoken language into text. In contrast, ALLMs aim to achieve a more holistic understanding of acoustic environments, encompassing not only the lexical content of speech but also paralinguistic cues (e.g., prosody, affective tone), speaker characteristics, musical elements, and background environmental sounds (Tang et al., 2024). Such deep exploration of the rich semantic information embedded in audio signals is crucial for realizing more natural and context-aware human-computer interaction. ALLMs are generally divided into two primary categories: speech understanding models and speech interaction models.
The rapid maturation of this field has been largely propelled by significant advancements in self-supervised learning (SSL) methodologies, which enable models to acquire robust representations from vast quantities of unlabeled audio data. Concurrently, sophisticated multimodal training paradigms have played a critical role, facilitating the synergistic integration and joint learning of information across auditory and linguistic modalities (Girdhar et al., 2023; Rubenstein et al., 2023; Zeng et al., 2024; Wang et al., 2024a). By aligning the acoustic feature space with the inherent semantic comprehension capabilities of LLMs, ALLMs are able to address tasks beyond simple speech-to-text conversion, such as audio event classification, audio scene description, audio-based question answering, and even engaging in multi-turn spoken dialogues. These capabilities mark new frontiers for developing artificial intelligence applications that can more profoundly interpret and respond to our auditory world. However, as ALLMs are increasingly integrated into real-world applications, understanding their impact under various trustworthiness conditions becomes critically important. This study aims to construct a benchmark, AudioTrust, to comprehensively and systematically evaluate the performance and potential risks of ALLMs across different trustworthiness dimensions, such as robustness, fairness, privacy protection, and safety. This evaluation is intended to provide scientific evidence and practical guidance for the responsible development, deployment, and regulation of ALLMs.
@@ A.1 Speech Understanding Models
Speech understanding models process and comprehend audio inputs, transforming them into semantic representations that facilitate language understanding. However, they lack the ability to generate audio responses. These models typically operate in a unidirectional manner, receiving audio as input and producing text-based outputs. Notable representatives include Qwen2-Audio (Chu et al., 2024), which integrates audio understanding capabilities into the Qwen2 (Yang et al., 2024a) via dedicated audio encoders and cross-modal adapters. These models demonstrate strong performance in tasks such as speech transcription, audio description, and audio-based question answering, yet their outputs remain restricted to textual modalities. SALMONN (Tang et al., 2024) likewise exhibits robust semantic audio understanding across diverse acoustic conditions, while maintaining a purely text-based output interface.
@@ A.2 Speech Interaction Models
Speech interaction models go beyond mere comprehension to enable bidirectional audio communication. These models are capable not only of understanding audio inputs, but also of generating contextually appropriate audio responses, thereby facilitating more natural human-computer interaction. Prominent examples include GPT-4o (OpenAI et al., 2024), which represents a significant advance in multimodal interactive capability by processing and generating audio in near real-time conversational scenarios. MiniCPM-o 2.6 (Yao et al., 2024) provides similar functionalities in an open-source format, supporting coherent audio dialogues while demonstrating comprehension of audio contexts. Such models enable a wide range of applications, from virtual assistants to assistive tools for visually impaired users.
@@ Appendix B Benchmark Models
To systematically investigate these trustworthiness aspects, we have selected a diverse set of models. This set includes both mainstream proprietary commercial models, such as GPT-4 (OpenAI et al., 2024) and Gemini (Team et al., 2023), as well as representative and robust open-source ALLMs, including Qwen2-Audio (Chu et al., 2024) and MiniCPM-o 2.6 (Yao et al., 2024). To ensure fairness and objectivity, all models are systematically tested on the same datasets and with identical evaluation metrics, followed by thorough comparative analyses of experimental results. It is worth noting that our methodology considers not only the fundamental audio comprehension capabilities of each model, but also examines their potential strengths and limitations in aspects such as complex interactions and knowledge transfer. This systematic safety evaluation provides a solid foundation for the future optimization and development of ALLMs.
@@ B.1 Open-Source Models
In conducting trustworthiness evaluations of unified ALLMs, we selected nine representative open-source audio and multimodal models: SALMONN, Ultravox, Qwen2-Audio, MiniCPM-o 2.6, Step-Fun, Qwen2.5-omni, Kimi-Audio, OpenS2S, Step-Audio2.
1. SALMONN (Tang et al., 2024) pioneered a dual-encoder architecture (Whisper speech encoder and BEATs audio encoder) together with a window-level Q-Former and LoRA adapters. This enables the pretrained Vicuna text LLM to achieve unified understanding of speech, environmental sounds, and music. The model also demonstrates emergent capabilities in cross-modal reasoning beyond the training tasks and in few-shot activation tuning.
2. Ultravox (AI, 2024) directly maps raw audio into the high-dimensional representation space of LLMs, thereby seamlessly eliminating the traditional ASR stage. This model not only comprehends speech content but also captures paralinguistic features such as tone and pauses, and supports streaming text outputs.
3. Qwen2-Audio (Chu et al., 2024) is a large-scale audio-language model that establishes a seamless pipeline between the Whisper-large-v3 encoder and the Qwen-7B language model, thereby supporting both spoken dialogue and audio analysis interaction modes. In real conversational and multitask zero-shot evaluations, the model leverages Mel-spectrograms of 16kHz audio combined with instruction tuning and Direct Preference Optimization (DPO), significantly improving the precision and robustness of responses to human intent.
4. MiniCPM-o 2.6 (Yao et al., 2024) integrates four major components: SigLip-400M, Whisper-medium, ChatTTS-200M, and Qwen2.5-7B, supporting bilingual real-time dialogue in an end-to-end multimodal fashion, as well as controllable interactions in emotion and speaking rate, and high-quality voice cloning. It consistently outperforms proprietary models of equivalent scale on benchmarks such as OpenCompass and StreamingBench.
5. Step-Fun (Huang et al., 2025) is a production-ready open-source real-time speech–text multimodal system that tackles data-collection cost, weak dynamic control, and limited intelligence via four pillars: a 130B unified understanding–generation model, a generative speech data engine enabling affordable voice cloning and distilling the lightweight Step-Audio-TTS-3B, an instruction-driven fine-control mechanism spanning dialects, emotions, singing, and rap, and an enhanced cognitive layer with tool calling and role-playing for complex tasks.
6. Qwen2.5-Omni (Xu et al., 2025b) builds upon Qwen2.5-VL/Audio by introducing the Thinker-Talker architecture and TMRoPE (Time-aligned Multimodal RoPE) temporal alignment embedding. This allows the model to stream and process text, image, audio, and video inputs concurrently within a single framework, with the ability to produce both textual and natural speech outputs in synchronization.
7. Kimi-Audio (Ding et al., 2025) is an open-source audio foundation model for understanding, generation, and conversation; it adopts a 12.5Hz audio tokenizer and an LLM-based architecture that ingests continuous features and emits discrete tokens, alongside a chunk-wise streaming detokenizer via flow matching for low-latency inference.
8. OpenS2S (Wang et al., 2025a) built on the BLSP-Emo empathetic speech-to-text backbone, it introduces a streaming interleaved decoding architecture for low-latency speech generation while capturing rich paralinguistic cues for expressive responses.
9. Step-Audio 2 (Wu et al., 2025) is an end-to-end multimodal LLM for industry-grade audio understanding and speech conversation, combining a latent audio encoder with reasoning-centric RL to boost ASR and audio comprehension; by folding discrete audio-token generation into language modeling, it becomes highly responsive to paralinguistic cues in real time.